# Impact of phylogeny on the inference of functional sectors from protein sequence data

### v1

@@ Results
Figure 7: Impact of earliest mutation generation $G$ on eigenvector components. Violin plots of the absolute value of components of the key eigenvectors of the ICOD, covariance $C$ and SCA matrices are represented versus the earliest mutation generation $G$ at which the associated site first mutates in the phylogeny. Results are shown for data sets generated with $\mu=50$ (top panels) and $\mu=5$ (bottom panels). Other parameter values are the same as in Fig 3. Violin plots are obtained over 100 realisations of data generation.
@@
How does ICOD perform at predicting sites with large mutational effects on natural data? Does its robustness to phylogeny give it an advantage? To address this, we consider natural protein families with published Deep Mutational Scan (DMS) experimental data, which we consider as ground truth. We investigate 30 protein families [46] listed in table S1. For each of them, we rank sites by the magnitude of predicted mutational effects using ICOD, SCA, Mutual Information (MI, see [45]), and conservation. We compare these predictions to DMS data. Specifically, for each protein family, we start from the reference sequence mutated in the DMS experiment, and construct a multiple sequence alignment of homologs of this protein whose depth can be varied by selecting the neighbors of the reference sequence in Jukes-Cantor distance up to a given phylogenetic cutoff [45]. Next, for each method, we compute the top eigenvector on MSAs constructed using different phylogenetic cutoffs, and then we summed component by component the eigenvectors corresponding to these different cutoffs. We use the components of the resulting vector as predictors of mutational effects, and the DMS data as ground truth. Fig 8 shows that all methods are able to predict important sites, the most successful ones being conservation and ICOD. Indeed, the average of the symmetrized AUC over all families is 0.45 for conservation, 0.43 for ICOD, 0.35 for MI and 0.26 for SCA. Furthermore, we observe that better results tend to be obtained for DMS with bimodal distributions of fitness effects (see Fig S12 for examples). This suggests that when the data comprises sites with substantially stronger mutational effects than others for the function probed by the DMS (i.e. a functional sector), then this sector can be well identified by the methods considered here, in particular conservation and ICOD.
@@
Figure 8: Identifying functionally important sites in natural protein families. The symmetrized AUC for the prediction of sites with large mutational effects is computed on 30 protein families, using four different methods: ICOD, SCA, MI and Conservation, using Deep Mutational Scan (DMS) data as ground truth. For ICOD and MI, the average product correction (APC) [4] is applied to the matrix of interest (it was found to improve the average performance over families for these methods, but not for SCA). For ICOD, MI and SCA, the components of the eigenvector associated to the largest eigenvalue are employed to make predictions of mutational effects. Protein families are ordered by decreasing symmetrized AUC for ICOD. The mapping between protein family number and name is given in Table S1. The protein families shaded in grey have DMS data featuring a unimodal shape, the other ones have a bimodal shape.
@@
While conservation and ICOD yield the best identification of functionally important sites, they rely on entirely different signals. Indeed, ICOD was designed to focus primarily on correlations and not on conservation – recall that the diagonal of the inverse covariance matrix is set to zero in order to eliminate conservation. Thus, we expect these two methods to find different important sites. Table S1 shows the number of sites identified as functionally important by both ICOD and conservation (“Both”) or only by either method (“ICOD” or “Cons.”), out of those that are deemed important by the DMS, i.e. the $L_{S}$ sector sites found by binarising the DMS data and setting a cutoff (see Methods). We find that predictions can indeed differ. For instance, $\beta$-lactamase has 12 sites that are predicted by ICOD only and 8 sites by conservation only, although the two methods yield a similar symmetrized AUC. We examined the 3D structures of HRas and PDZ and the location of the sites identified by both methods or by only one method. In these two cases, DMS data is obtained by focusing on ability to bind other molecules. Accordingly, functionally important sites identified in these DMS are located around interaction regions. We observe that the sites detected by ICOD only tend to be closer to these interfaces than those detected by conservation only. These sites have large DMS scores and are localized close to binding interfaces, strongly hinting at their functional importance, while they are not among the most conserved ones. Interestingly, ICOD allows to reveal such sites in these examples.
@@ Discussion
Next, using synthetic data from a minimal model where we can tune the amounts of phylogeny and of nonlinear selection on an additive trait, we showed that phylogenetic correlations can substantially impair the inference of functional sectors from sequence data. However, ICOD and conservation are more robust to phylogenetic noise than SCA and covariance. The robustness of ICOD suggests that focusing on the small-eigenvalue modes of the covariance matrix is a successful strategy, as the large-eigenvalue ones tend to be most affected by phylogeny [35, 36]. Conservation has been used a lot to identify functionally important sites, and thus, its good performance is not surprising. In addition, it was argued in [48] that at least in some cases, the success of SCA can be attributed to its use of conservation. Here, we find that SCA’s performance is indeed often tied to that of conservation. Importantly, in ICOD, the diagonal terms of the inverse covariance matrix are set to zero in order to eliminate the impact of conservation as much as possible, and focus on correlations. Thus, the robustness of ICOD cannot be explained by that of conservation, and the two methods should provide different and complementary information. While SCA aimed to combine the two important and complementary ingredients of conservation and covariance, our analysis of controlled synthetic data suggests that considering the two separate ingredients and using ICOD instead of covariance to further separate them may yield even more information.
@@
Our analysis of several natural protein families for which DMS data is available showed that conservation and ICOD were the best predictors of mutational effects among the methods we considered. Our results regarding synthetic data suggest that this could be due to the robustness of these methods to phylogenetic correlations, which are bound to exist in natural sequence data. Here, we showed that phylogenetic correlations make the inference of functional sectors challenging, very much like they obscure the inference of structural contacts [4, 36, 37, 38, 39, 40]. It is important to note that phylogenetic correlations are nevertheless interesting and provide useful signal e.g. for the inference of protein partners among paralogs [49, 47, 50].
@@
Our model of selection on an additive trait with a quadratic Hamiltonian is formally close to Potts models [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23] and to generalized Hopfield models [51, 52]. Potts models have allowed mutational effect analysis [53, 46, 54, 55]. However, our selection model induces a specific dependence of the fields and of the couplings in the mutational effect vector which does not generically exist in these models. This is what makes it possible to recover mutational effects directly via an eigenvector of the ICOD matrix, instead of needing to compare inferred energies [54]. Generalizing beyond quadratic models would be very interesting [56, 57]. Besides, while we started from a generic model for sectors which involves an additive trait [33, 34, 32] under nonlinear natural selection, it would be interesting to analyze specific traits in more detail. One of them [31] is the energetic cost of protein deformations within an elastic-network model [58]. The low-energy deformation modes of protein structures are important in many functionally important protein deformations [59, 60, 61, 62, 63], and are robust to sequence variation within protein families [64, 65, 66]. It would be interesting to reexamine these cases with a more in-depth analysis of phylogenetic effects.
@@ Methods
In order to incorporate phylogeny, we take a functional sequence generated as above as the ancestor, and let it evolve on a binary branching tree, see Fig 1B. A fixed number $\mu$ of accepted mutations are performed independently along each branch, and the sequences sequences at the tree leaves constitute our MSA. Mutations are accepted with probability $p$ from Eq 2, which maintains the same selection on the trait. The resulting sequences contain correlations from phylogeny (controlled by $\mu$) in addition to those coming from selection. If $\mu$ is large enough, even sister sequences become independent and equilibrium statistics are recovered. Conversely, if $\mu$ is small, sister sequences are very similar and phylogeny is strong. Note that we can also generate sequences wit phylogeny and no selection using the same method, but accepting all proposed mutations.
@@
Deep Mutational Scan (DMS) experiments provide a direct measurement of the fitness effect associated to each possible point mutation at each site of a reference sequence. In DMS experiments, selection for a given function (e.g. ability to bind to another molecule) is applied on the mutated and the wild type sequences. The ratio of the number of sequences after and before selection is called $r_{WT}$ for the wild-type and $r_{mut}$ for the mutant. Most often, the fitness effect of the particular mutation is assessed via the logarithm of the enrichment ratio $\log\left(r_{mut}/r_{WT}\right)$. Functionally important sites are those for which mutations are most costly, leading to a small $r_{mut}$ and a negative log enrichment score. Thus, we construct the analogue of our vector $\vec{D}$ of mutational effects by taking for each site the most negative (smallest) enrichment ratio, associated with the most deleterious mutation at this site. We start from the 30 DMS datasets collected in [46]. We binarise the minimum enrichment scores from the DMS data. We observe that most score distributions are in practice either bimodal or unimodal, see Fig S12 for some examples. Thus, we fit for each family either a linear combination of two Gaussian distributions or a single Gaussian distribution to find a cutoff value for the binarisation. For the unimodal families, we use the mean of the fitted Gaussian distribution as the cutoff, while for the bimodal ones we use the position of the minimum between the two peaks. In some cases, there are more than two peaks in the distribution: then, we fit a bimodal distribution to the two most negative peaks, see upper middle panel in Fig S12. In all cases, we define the sector as comprising the sites with values more negative than the cutoff value. The number of thus-defined sector sites is called $L_{S}$.
@@
In order to infer sectors in natural data, we need to construct Multiple Sequence Alignments (MSAs) of homologs of the reference sequence of the DMS experiment. For this, we use the method from reference [46]. We search the reference sequence against the UniRef100 database using jackhmmer from the HMMER3 suite [71], specifically the command "jackhmmer --incdomT 0.5*L --cpu 6 -N 5 -A pathsave pathrefseq pathuniref", where $L$ is the length of the wild type sequence. We restrict to match states where the reference sequence does not have a gap, remove columns with more than 30% gaps, and remove sequences which have more than 20% gaps. Inspired by reference [45], we further subsample the MSA to generate a collection of MSAs comprising neighbors of the reference sequence up to different phylogenetic cutoffs. Specifically, we only retain sequences within a given Jukes-Cantor distance, termed cutoff distance, of the reference sequence. Each gap of the MSA is then replaced by the amino acid possessed at the same site by the nearest sequence in terms of Jukes-Cantor distance. We employ the following values of cutoff distance: 0.4, 0.6, 0.8, 0.9, 1.0, 1.1, 1.2, 1.3, 1.4, 1.5, 1.6, 1.7, 1.8, 1.9 and 2.0, yielding 15 MSAs for each protein family. Note that for protein families 1, 3, and 27, MSAs were generated only up to cutoffs 1.2, 1.4 and 1.4 respectively, due to computational constraints.
@@
In the case of synthetic data, we apply each method (ICOD, covariance, SCA, conservation) to the generated MSA. In the case of natural data, we have 15 different MSAs for each protein family. We apply each method (ICOD, MI, SCA, conservation) on each MSA and extract the eigenvector associated to the eigenvalue of focus. This yields 15 different eigenvectors that are added together component by component. Because normalized eigenvectors are defined up to an overall sign, we set the sign of one eigenvector arbitrarily and choose the one of the other eigenvectors such that they have a positive Pearson correlation with the first eigenvector. The overall sign of the sum remains arbitrary, but we employ scoring methods that are invariant to it.
@@
For natural data, a similar construction can be made by employing the covariance matrix in Eq 7 with pseudocount-corrected frequencies (Eq 3 and 4), with a pseudocount value $a=0.05$. The reference-sequence gauge [31] is used, and we do not compute the frequency of the amino acid in the reference sequence at a given position, yielding a covariance matrix of size $(19\times L,19\times L)$. This matrix is then inverted, the Frobenius norm is taken on each submatrix of size $19\times 19$ corresponding to each pair of sites $(i,j)$. Next, the diagonal is set to zero.
@@
For natural data, we mainly use the symmetrized AUC to evaluate performance, because DMS only test one aspect of function and the measured fitness can involve additional nonlinearities not captured by our simple model.

### v2

@@ Results
Figure 5: Impact of earliest mutation generation $G$ on eigenvector components. Violin plots of the absolute value of components of the key eigenvectors of the ICOD, covariance $C$ and SCA matrices are represented versus the earliest mutation generation $G$ at which the associated site first mutates in the phylogeny. Results are shown for data sets generated with $\mu=50$ (top panels) and $\mu=5$ (bottom panels). Datasets of $M=2048$ sequences of length $L=200$ were generated along a perfect binary tree with 11 generations, using two different numbers $\mu$ of accepted mutations per branch. As in Fig 3, we employed the mutation acceptance criterion in Eq 2 with $\kappa=10/\left(\sum_{i}{D_{i}^{2}}\right)$ and $\tau^{*}=90$. We used the same vector of mutational effect $\vec{D}$ as in Fig 2 and Fig 3. Violin plots are obtained over 100 realisations of data generation.
@@
How does ICOD perform at predicting sites with large mutational effects on natural data? Does its robustness to phylogeny give it an advantage? To address this, we consider natural protein families with published Deep Mutational Scan (DMS) experimental data, which we consider as ground truth. We investigate 30 protein families [46] listed in table S1. For each of them, we rank sites by the magnitude of predicted mutational effects using ICOD, SCA, Mutual Information (MI, see [45]), and conservation. We compare these predictions to DMS data. Specifically, for each protein family, we start from the reference sequence mutated in the DMS experiment, and construct a multiple sequence alignment of homologs of this protein whose depth can be varied by selecting the neighbors of the reference sequence in Jukes-Cantor distance up to a given phylogenetic cutoff [45]. Next, for each method, we compute the top eigenvector on MSAs constructed using different phylogenetic cutoffs, and then we summed component by component the eigenvectors corresponding to these different cutoffs. We use the components of the resulting vector as predictors of mutational effects, and the DMS data as ground truth. Note that there are other ways of aggregating scores across phylogenetic cutoffs. Averaging performance metrics across cutoffs gives results that are overall consistent with those obtained by summing eigenvector components, but slightly less good, see Fig S8. Fig 6 shows that all methods are able to predict important sites, the most successful ones being conservation and ICOD. Indeed, the average of the symmetrized AUC over all families is 0.45 for conservation, 0.43 for ICOD, 0.35 for MI and 0.26 for SCA. Furthermore, we observe that better results tend to be obtained for DMS with bimodal distributions of fitness effects (see Fig S9 for examples). This suggests that when the data comprises sites with substantially stronger mutational effects than others for the function probed by the DMS (i.e. a functional sector), then this sector can be well identified by the methods considered here, in particular conservation and ICOD.
@@
Figure 6: Identifying functionally important sites in natural protein families. The symmetrized AUC for the prediction of sites with large mutational effects is computed on 30 protein families, using four different methods: ICOD, SCA, MI and Conservation, using Deep Mutational Scan (DMS) data as ground truth. For ICOD and MI, the average product correction (APC) [4] is applied to the matrix of interest (it was found to improve the average performance for most families for these methods, but not for SCA). For ICOD, MI and SCA, the components of the eigenvector associated to the largest eigenvalue are employed to make predictions of mutational effects. Protein families are ordered by decreasing symmetrized AUC for ICOD. The mapping between protein family number and name is given in Table S1. The protein families shaded in grey have DMS data featuring a unimodal shape, the other ones have a bimodal shape.
The observation that ICOD and conservation tend to outperform SCA is consistent with our results obtained in synthetic data. To assess more directly the link between these results and phylogeny in natural data, Fig S10 shows the relative difference of ICOD and SCA performances versus MSA diversity, characterized by the mean pairwise Hamming distance between sequences in the MSA (see also Table S2). We find that for all 7 families with mean pairwise Hamming distance smaller than 0.55, ICOD significantly outperforms SCA, while results are more mixed for more diverse families. These 7 least diverse families are precisely those for which we expect the effect of phylogeny to be strongest. This result suggests that phylogeny does explain part of the success of ICOD on these natural datasets.
@@
While conservation and ICOD yield the best identification of functionally important sites, they rely on entirely different signals. Indeed, ICOD was designed to focus primarily on correlations and not on conservation – recall that the diagonal of the inverse covariance matrix is set to zero in order to eliminate conservation. Accordingly, with synthetic data, we observe in Fig II in section S2 of the Supplementary Appendix that ICOD reaches very good performance for a favored trait value $\tau^{*}=0$, such that conservation is uninformative. Thus, we expect these two methods to find different important sites. Table S1 shows the number of sites identified as functionally important by both ICOD and conservation (“Both”) or only by either method (“ICOD” or “Cons.”), out of those that are deemed important by the DMS, i.e. the $L_{S}$ sector sites found by binarising the DMS data and setting a cutoff (see Methods). There is often substantial overlap between sites found by ICOD and by conservation, which suggests that in practice many important sites are both somewhat conserved and somewhat correlated with others. Nevertheless, we find that the predictions of ICOD and of conservation can substantially differ, in line with our expectations. For instance, $\beta$-lactamase has 12 sites that are predicted by ICOD only and 8 sites by conservation only, although the two methods yield a similar symmetrized AUC. We examined the 3D structures of HRas and PDZ and the location of the sites identified by both methods or by only one method in Fig S12 and S13. In these two cases, DMS data is obtained by focusing on ability to bind other molecules. Accordingly, functionally important sites identified in these DMS are located around interaction regions. We observe that the sites detected by ICOD only tend to be closer to these interfaces than those detected by conservation only. These sites have large DMS scores and are localized close to binding interfaces, strongly hinting at their functional importance, while they are not among the most conserved ones. Interestingly, ICOD allows to reveal such sites in these examples.
@@ Discussion
Next, using synthetic data from a minimal model where we can tune the amounts of phylogeny and of nonlinear selection on an additive trait, we showed that phylogenetic correlations can substantially impair the inference of functional sectors from sequence data. However, ICOD and conservation are generally more robust to phylogenetic noise than SCA and covariance. The robustness of ICOD suggests that focusing on the small-eigenvalue modes of the covariance matrix is a successful strategy, as the large-eigenvalue ones tend to be most affected by phylogeny [35, 36]. Conservation has been used a lot to identify functionally important sites (see e.g. [48]), and thus, its good performance is not surprising. In addition, it was argued in [50] that at least in some cases, the success of SCA can be attributed to its use of conservation. Here, we find that SCA’s performance is indeed often tied to that of conservation. Importantly, in ICOD, the diagonal terms of the inverse covariance matrix are set to zero in order to eliminate the impact of conservation as much as possible, and focus on correlations. Thus, the robustness of ICOD cannot be explained by that of conservation, and the two methods should provide different and complementary information. While SCA aimed to combine the two important and complementary ingredients of conservation and covariance, our analysis of controlled synthetic data suggests that considering the two separate ingredients and using ICOD instead of covariance to further separate them may yield even more information.
@@
Our analysis of several natural protein families for which DMS data is available showed that conservation and ICOD were the best predictors of mutational effects among the methods we considered. This could be due to the robustness of these methods to phylogenetic correlations, especially given that ICOD particularly outperforms SCA for natural protein families with small diversity, where phylogenetic correlations should be most important. Here, we showed that phylogenetic correlations make the inference of functional sectors challenging, very much like they obscure the inference of structural contacts [4, 36, 37, 38, 39, 40]. It is important to note that phylogenetic correlations are nevertheless interesting and provide useful signal e.g. for the inference of protein partners among paralogs [51, 47, 52].
@@
Our model of selection on an additive trait with a quadratic Hamiltonian is formally close to Potts models [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23] and to generalized Hopfield models [53, 54]. Potts models have allowed mutational effect analysis [55, 46, 56, 57]. However, our selection model induces a specific dependence of the fields and of the couplings in the mutational effect vector which does not generically exist in these models. This is what makes it possible to recover mutational effects directly via an eigenvector of the ICOD matrix, instead of needing to compare inferred energies [56]. Methods able to predict mutational effects can be used to determine sites with top mutational effects. We compared using the ICOD eigenvector and using the Potts model energy for this as in [46], and found that the latter had slightly better average symmetrized AUC. Note however that the results of [46] were obtained using different MSAs, which makes the comparison imperfect. We stress that eigenvector-based methods, such as SCA and ICOD, have the potential to disentangle different aspects of function via different eigenvectors [27, 30, 31], which is not the case of Potts energy-based methods. Studying the impact of phylogeny on cases with simultaneous selection on multiple traits, as in Ref. [31], would be interesting. Furthermore, generalizing beyond quadratic models would be very interesting [58, 59]. Besides, while we started from a generic model for sectors which involves an additive trait [33, 34, 32] under nonlinear natural selection, it would be interesting to analyze specific traits in more detail. One of them [31] is the energetic cost of protein deformations within an elastic-network model [60]. The low-energy deformation modes of protein structures are important in many functionally important protein deformations [61, 62, 63, 64, 65], and are robust to sequence variation within protein families [66, 67, 68]. It would be interesting to reexamine these cases with a more in-depth analysis of phylogenetic effects.
@@ Methods
In order to incorporate phylogeny, we take a functional sequence generated as above as the ancestor, and let it evolve on a perfect binary tree, see Fig 1B. A fixed number $\mu$ of accepted mutations are performed independently along each branch, and the sequences sequences at the tree leaves constitute our MSA. Mutations are accepted with probability $p$ from Eq 2, which maintains the same selection on the trait. The resulting sequences contain correlations from phylogeny (controlled by $\mu$) in addition to those coming from selection. If $\mu$ is large enough, even sister sequences become independent and equilibrium statistics are recovered. Conversely, if $\mu$ is small, sister sequences are very similar and phylogeny is strong. Note that we can also generate sequences with phylogeny and no selection using the same method, but accepting all proposed mutations.
@@
Deep Mutational Scan (DMS) experiments provide a direct measurement of the fitness effect associated to each possible point mutation at each site of a reference sequence. In DMS experiments, selection for a given function (e.g. ability to bind to another molecule) is applied on the mutated and the wild type sequences. The ratio of the number of sequences after and before selection is called $r_{WT}$ for the wild-type and $r_{mut}$ for the mutant. Most often, the fitness effect of the particular mutation is assessed via the logarithm of the enrichment ratio $\log\left(r_{mut}/r_{WT}\right)$. Functionally important sites are those for which mutations are most costly, leading to a small $r_{mut}$ and a negative log enrichment score. Thus, we construct the analogue of our vector $\vec{D}$ of mutational effects by taking for each site the most negative (smallest) enrichment ratio, associated with the most deleterious mutation at this site. We start from the 30 DMS datasets collected in [46], listed in Table S1, and with metrics describing their diversity and depth given in Table S2. To define the set of sites with large mutational effects (which corresponds to the sector), we use a threshold on the DMS measurements to binarise them, leading sector and non-sector sites. The threshold is set as follows. We observe that most score distributions are in practice either bimodal or unimodal, see Fig S9 for some examples. Thus, we fit for each family either a linear combination of two Gaussian distributions or a single Gaussian distribution to find a cutoff value for the binarisation. For the unimodal families, we use the mean of the fitted Gaussian distribution as the cutoff, while for the bimodal ones we use the position of the minimum between the two peaks. In some cases, there are more than two peaks in the distribution: then, we fit a bimodal distribution to the two most negative peaks, see upper middle panel in Fig S9. In all cases, we define the sector as comprising the sites with values more negative than the cutoff value. The number of thus-defined sector sites is called $L_{S}$. Once the sector and non-sector sites are defined, the performance of various models at identifying sector sites is quantified by the symmetrized AUC. See the paragraph on “Performance evaluation” below for details.
@@
In order to infer sectors in natural data, we need to construct Multiple Sequence Alignments (MSAs) of homologs of the reference sequence of the DMS experiment. For this, we use the method from reference [46]. We search the reference sequence against the UniRef100 database using jackhmmer from the HMMER3 suite [73], specifically the command "jackhmmer --incdomT 0.5*L --cpu 6 -N 5 -A pathsave pathrefseq pathuniref", where $L$ is the length of the wild type sequence. We restrict to match states where the reference sequence does not have a gap, remove columns with more than 30% gaps, and remove sequences which have more than 20% gaps. Inspired by reference [45], we further subsample the MSA to generate a collection of MSAs comprising neighbors of the reference sequence up to different phylogenetic cutoffs [45]. Specifically, we only retain sequences up to a given maximum Jukes-Cantor distance [74], which we refer to as “phylogenetic cutoff”, of the reference sequence. Each gap of the MSA is then replaced by the amino acid possessed at the same site by the nearest sequence in terms of Jukes-Cantor distance. We employ the following values of cutoff distance: 0.4, 0.6, 0.8, 0.9, 1.0, 1.1, 1.2, 1.3, 1.4, 1.5, 1.6, 1.7, 1.8, 1.9 and 2.0, yielding 15 MSAs for each protein family. Note that for protein families 1, 3, and 27, MSAs were generated only up to cutoffs 1.2, 1.4 and 1.4 respectively, due to computational constraints.
@@
In the case of synthetic data, we apply each method (ICOD, covariance, SCA, conservation) to the generated MSA. In the case of natural data, we have 15 different MSAs for each protein family. We apply each method (ICOD, MI, SCA, conservation) on each MSA and extract the eigenvector associated to the eigenvalue of focus. This yields 15 different eigenvectors that are added together component by component. Because normalized eigenvectors are defined up to an overall sign, we set the sign of one eigenvector arbitrarily and choose the one of the other eigenvectors such that they have a positive Pearson correlation with the first eigenvector. The overall sign of the sum remains arbitrary, but we employ scoring methods that are invariant to it. Note that for natural data, we do not aim to infer mutational effect values but only their rank, in line with the usual practice of SCA [27, 29]. While in principle we can infer mutational effect values from eigenvectors at least with ICOD [31], the measured fitness can involve additional nonlinearities not captured by our simple model.
@@
For natural data, a similar construction can be made by employing the covariance matrix in Eq 7 with pseudocount-corrected frequencies (Eq 3 and 4), with a pseudocount value $a=0.05$. The reference-sequence gauge [31] is used, meaning that the reference sequence of the dataset (i.e., the wild-type sequence for deep mutational scan data) is chosen as our baseline. Here, in practice, we do not compute the frequency of the amino acid in the reference sequence at a given position, yielding a covariance matrix of size $(19\times L,19\times L)$. This matrix is then inverted, the Frobenius norm is taken on each submatrix of size $19\times 19$ corresponding to each pair of sites $(i,j)$. Next, the diagonal is set to zero.
@@
For natural data, we mainly use the symmetrized AUC to evaluate performance, because DMS only test one aspect of function and the measured fitness can involve additional nonlinearities not captured by our simple model. To calculate the AUC, we vary the threshold on the numbers of sites predicted (which are the top eigenvector components). Specifically, computing the true positive and false positive rates for each threshold value allows constructing the receiver operating characteristic curve and computing the AUC.